OpenStudy (anonymous):

Chris is making lemonade. The directions say to mix 5 cups of water with 2 scoops of powdered lemonade mix. How many cups of water should he use with 2.5 scoops of fruit mix?

OpenStudy (anonymous):

proportion time 5/2 = x/2.5 (5 cups water to 2 scoops lemonade = x cups to 2.5 scoops) cross multiply (2)(x) = (5)(2.5) 2x = 12.5 x = 12.5/2 x = 6.25 or 6 1/4 scoops you can even check your answers.. 5/2 = 6.25/2.5 cross multiply (2)(6.25) = (5)(2.5) 12.5 = 12.5 (correct) -- if they come out equal, like this, then the answer is correct